OpenStudy (anonymous):

Given the function f(x) = 5x - 9, find f(-3).

OpenStudy (kaederfds):

f(-3) = 5.-3 - 9 = -15 -9 = -24

OpenStudy (anonymous):

-24 is the final answer ? & Thank you

OpenStudy (anonymous):

-24

OpenStudy (kaederfds):

yes, you're welcome